# Faster cosmological analysis with power spectrum without simulations

### 31 Jan 2024 v2

@@ 4.1.1 Application to data
We test the procedure’s suitability described above by constructing individual analytical covariance matrices for each of the 9 clustering samples we consider. We then treat the full analytical covariance matrix as a block diagonal matrix with the individual covariance matrix sitting in its corresponding block. We assume there is no correlation between different surveys mainly because of the large cosmological distances and weak overlap between the different samples. However within the BOSS survey, there is overlap between the $z1$ and $z2$ bins which cannot be modelled without extending the analytical covariance matrix work of Wadekar et al. (2020). For simplicity, we instead first test the effect of neglecting these cross-correlation terms entirely.
We fit our sample with the full simulated covariance matrix and a block-diagonal simulated covariance matrix where we remove the cross-correlation between the BOSS samples, replacing all non-block-diagonal terms with zero. The priors for our fit are summarized in Table 1, and match those in Glanville et al. (2022) where we fit the same data set with PyBird. The results are summarized in Fig 2 and Table 2. We found negigible shifts in the mean of the posteriors before and after removing the cross-correlation. There was a small change of <10% in the parameter uncertainties for $A_{s}$ and $h$, however this difference is very small and it is not clear if this is due to sampling noise, because of the absence of physical correlation between these two bins, or actually because the simulated cross-correlation itself is noisy. As such, we deem it reasonable to simply ignore this cross-correlation component in our fits, still making sure to remove it from both the simulated and analytic covariances for any comparisons are on equal footing.
@@
Table 1: The priors on the cosmological parameters and the nuisance parameters for this work. Here, $\mathcal{U}$($\mathcal{N}$) denotes the uniform (Gaussian) prior where the first number is the lower bound (mean) and the second number is the upper bound (standard deviation). These priors are used and validated by Glanville et al. (2022) to fit the same data set. $A_{s}^{\mathrm{fid}},h^{\mathrm{fid}},$ and $(\Omega_{\mathrm{cdm}}h^{2})^{\mathrm{fid}}$ denotes the fiducial cosmological parameters used to calculate the center of the pre-computed grid of power spectrum. The widths of the priors (0.8, 0.08, and 0.04 for $A_{s}^{\mathrm{fid}},h^{\mathrm{fid}},$ and $(\Omega_{\mathrm{cdm}}h^{2})^{\mathrm{fid}}$ respectively) are to ensure cosmological parameters outside the grid won’t be evaluated. The fiducial parameter values can be found in Table 3, but are changed depending on the exact tests we run.
Figure 2: This plot compares the constraints from the Full simulated covariance matrix and the block-diagonal simulated covariance matrix. Table 2 summarizes the constraints of the cosmological parameters. In general, we see no significant changes in the cosmological parameters, which indicates ignoring the cross-correlation does not affect the mean of the posterior for our data-set.
@@
Figure 3: This plot compares the analytical covariance matrix with the simulated one for the eBOSS LRG NGC sample. The analytical covariance matrix is generated using the best-fit cosmological and bias parameters from a first fitting iteration, which itself used the “Template" parameters in Table 3 and no MOPED compression (A1 in Table 4). The blue band corresponds to the sampling noise of the simulated covariance matrix. From left to right, we show the auto-covariance matrix for the monopole, the cross-covariance matrix for the monopole and quadrupole, and the auto-covariance matrix for the quadrupole. From top to bottom, we fix the first wavevector of the covariance matrix to $0.045h\rm{Mpc}^{-1}$, $0.095h\rm{Mpc}^{-1}$, $0.145h\rm{Mpc}^{-1}$, and $0.195h\rm{Mpc}^{-1}$ and vary the second wavevector from $0.0h\rm{Mpc}^{-1}$ to $0.30h\rm{Mpc}^{-1}$. The plots show that the analytical auto-covariance matrix is generally consistent with the simulated covariance matrix within the sampling noise limit, especially along the diagonal. The small disagreement along the off-diagonal does not affect the cosmological constraints as shown in section 5.
Fig. 3 compares the analytical covariance matrix with the simulated covariance matrix for the eBOSS LRG NGC subsample. For the analytic covariance matrix we use the best-fit cosmological parameters obtained from our first fitting iteration which itself used an analytical covariance matrix computed with the “Template" parameters in Table 3, and without the compression (A1 in Table 4). [^5] You can find the comparison of the covariance matrix for the other 8 samples here https://github.com/YanxiangL/Analytical-compressed-cov/tree/main/analytic_vs_sim. The blue band is the sampling noise of the simulated covariance matrix computed with equation (4). Fig 3 illustrates that the analytical covariance matrix is generally consistent with the simulated covariance matrix within the sampling noise, especially along the diagonal. The small disagreement could be because the local Lagrangian relation is not perfect, so some part of the analytical covariance matrix deviates slightly from the simulated one. Nonetheless, we find that the disagreement in the off-diagonal elements of the covariance matrix does not significantly change the constraints on the cosmological parameters, as the amplitude of these components is generally much smaller than either the diagonals of their respective blocks. Wadekar et al. (2020) similarly demonstrates that the off-diagonal components of the covariance between the monopole and quadrupole have a marginal impact on the constraint of the cosmological parameters.
@@ 4.2 The MOPED algorithm
The Massively Optimized Parameter Estimation and Data compression (MOPED) algorithm has been demonstrated to be able to reduce the number of data points without losing information (Heavens et al., 2000; Heavens et al., 2017). The algorithm achieves this by keeping the Fisher Matrix of the statistic in question with respect to the parameters of interest the same before and after the compression. Section 2 demonstrates the compressed precision matrix gives a smaller fractional error and bias than the uncompressed precision matrix. Therefore, the compression could also give more reliable constraints for the cosmological parameters. Here we demonstrate its application to the EFTofLSS model and our large-scale structure dataset.
@@
where the model power spectrum is given by $\boldsymbol{P}$ and $\boldsymbol{P}_{,m}$ is the partial derivative of the model power spectrum with the $m^{\mathrm{th}}$ free parameter. The transpose is denoted by $T$ in equation (15).
To compute the compression matrix $\operatorname{\boldsymbol{\mathsf{D}}}$, we need to know the derivative of the model power spectrum with respect to the cosmological and nuisance parameters. For the derivatives with respect to the cosmological parameters, this is done when calculating the Taylor expansion of the power spectrum. A detailed description is given in section 4.3 and in this section, we mainly focus on the derivatives of the power spectrum with respect to the nuisance parameters. In PyBird, the non-linear power spectrum calculation is separated into 24 different terms (3 linear terms, 12 loop terms, 6 counter terms, and 3 stochastic terms). Each of these terms is independent of the nuisance parameters. More importantly, the final non-linear power spectrum is the linear combination of these 24 different terms multiplied by the appropriate bias parameters
@@
The MOPED compression changes the covariance matrix and the power spectra, so we have to change the likelihood function accordingly. Before applying the compression, the power spectrum is expected to follow a Gaussian distribution, so the likelihood function $L$ is given by
@@
where $\boldsymbol{P}_{d}$ is the data power spectrum and $\boldsymbol{P}^{W}=\operatorname{\boldsymbol{\mathsf{W}}}\boldsymbol{P_{l}}$ is the model power spectrum after convolving with the window function $\operatorname{\boldsymbol{\mathsf{W}}}$. In this work, both the analytical and simulated covariance matrix take the survey window function into account. Therefore, to calculate the compression matrix $\operatorname{\boldsymbol{\mathsf{D}}}$, we need to input the derivative of the non-linear power spectrum with the window function ($\boldsymbol{P}^{W}$) with respect to the cosmological and nuisance parameters. Furthermore, this means the compressed windowed non-linear power spectrum is given by the compression matrix $\operatorname{\boldsymbol{\mathsf{D}}}$ multiplied by the windowed nonlinear power spectrum ($\boldsymbol{P}^{W}$). From Heavens et al. (2000), the covariance matrix after compression is the identity matrix, so the likelihood after applying MOPED is given by
@@
[…]
@@
$F_{2,ij}=(P^{W}_{,b_{G_{i}}})^{T}\operatorname{\boldsymbol{\Psi}}P^{W}_{,b_{G_{j}}}$
$F_{1,i}=(P^{W}_{\mathrm{const}})^{T}\operatorname{\boldsymbol{\Psi}}P^{W}_{,b_{G_{i}}}+P_{d}^{T}\operatorname{\boldsymbol{\Psi}}P^{W}_{,b_{G_{i}}}$
$F_{0}=-\frac{1}{2}(P^{W}_{\mathrm{const}})^{T}\operatorname{\boldsymbol{\Psi}}P^{W}_{\mathrm{const}}+(P^{W}_{\mathrm{const}})^{T}\operatorname{\boldsymbol{\Psi}}P_{d}-\frac{1}{2}P_{d}^{T}\operatorname{\boldsymbol{\Psi}}P_{d},$ (20)
and
$P^{W}_{,b_{G_{i}}}$ $=\frac{\partial P^{W}}{\partial b_{G_{i}}}\biggl{|}_{\vec{b_{G}}\rightarrow 0},$
$P^{W}_{\mathrm{const}}$ $=P^{W}\bigl{|}_{\vec{b_{G}}\rightarrow 0},$ (21)
with $\vec{b_{G}}=\{b_{3}^{\mathrm{EFT}},c_{ct},c_{r,1},c_{r,2},c_{\epsilon,1},c_{\epsilon,\mathrm{mono}},c_{\epsilon,\mathrm{quad}}\}$.
Since now the marginalized likelihood depends on $P^{W}_{,b_{G_{i}}}$ and $P^{W}_{\mathrm{const}}$ instead of $\boldsymbol{P}^{W}$, we want to know how to apply the compression to the marginalized likelihood.
d'Amico et al. (2020) shows
$P^{W}=\sum_{i}b_{G,i}P^{W}_{,b_{G_{i}}}+P_{\mathrm{const}},.$ (22)
@@
Therefore, the windowed nonlinear power spectrum $\boldsymbol{P}^{W}$ is just a linear combination of $P^{W}_{,b_{G_{i}}}$ and $P^{W}_{\mathrm{const}}$. The bias parameters are just constants and would not affect the matrix multiplication, so we have
$DP^{W}=\sum_{i}b_{G,i}DP^{W}_{,b_{G_{i}}}+DP_{\mathrm{const}}.$ (23)
@@
This means we can calculate the compression matrix for $P^{W}_{,b_{G_{i}}}$ and $P^{W}_{\mathrm{const}}$ the same way as $P^{W}$, by multiplying the compression matrix to it. After some mathematical manipulation, the expression for the marginalized likelihood stays the same, but one needs to change equation (20) to
$F_{2,ij}^{\mathrm{comp}}=(DP^{W}_{,b_{G_{j}}})^{T}DP^{W}_{,b_{G_{i}}}$
$F_{1,i}^{\mathrm{comp}}=(DP^{W}_{,b_{G_{i}}})^{T}DP^{W}_{\mathrm{const}}+(DP^{W}_{,b_{G_{i}}})^{T}DP_{d}$
$F_{0}^{\mathrm{comp}}=-\frac{1}{2}(DP^{W}_{\mathrm{const}})^{T}DP^{W}_{\mathrm{const}}+(DP_{d})^{T}DP^{W}_{\mathrm{const}}-\frac{1}{2}(DP_{d})^{T}DP_{d}.$ (24)
@@
There are many ways to calculate the equation (24) by rearranging the order of the matrix/vector multiplications. For example, for $F_{2,ij}^{\mathrm{comp}}$ we can write it as $(P^{W}_{,b_{G_{i}}})^{T}D^{T}DP^{W}_{,b_{G_{j}}}$. However, it is not computationally efficient to calculate $F_{2,ij}^{\mathrm{comp}}$ this way because $D^{T}D$ has the same dimension as the uncompressed precision matrix, which will not speed up the code. Having investigated different approaches, we found the most computationally efficient way was to pre-compute $DP_{d}$ and $(DP_{d})^{T}DP_{d}$ because the data vector does not change during MCMC. Then during each step in MCMC, we calculate $DP^{W}_{,b_{G_{i}}}$ and $DP^{W}_{\mathrm{const}}$ first, then we use the transpose operation to find $(DP^{W}_{,b_{G_{i}}})^{T}$ and $(DP^{W}_{\mathrm{const}})^{T}$ because the transpose operation is much faster than matrix multiplication. After transposing, both the model and data power spectrum are compressed, so now the matrix multiplication is faster. We then substitute the compressed power spectra into equation (24) and equation (19) to find the marginalized likelihood.
@@
Compression has several benefits. The MOPED algorithm significantly reduces the dimension of the data vector. Therefore, to reach the same precision, we will need fewer simulations to estimate the covariance matrix as shown in Fig 1. Additionally, the chi-squared calculation is one of the slowest parts when sampling the posterior with MCMC. Reducing the size of the data vector speeds up the MCMC algorithm and we don’t have to multiply by the covariance matrix after the compression because any matrix multiplied by the identity matrix is itself.
@@
However, there are also caveats — we need to multiply the compression matrix with the model power spectrum at every step. This could cancel out the potential speed-up. Therefore, we will also test whether the MOPED compression is able to speed up the posterior sampling. The second caveat is, theoretically, the compression matrix needs to be generated with the best-fit parameters for the compression to be lossless. Heavens et al. (2000) found that using parameters that are different from the best-fit only marginally increases the credible regions. We demonstrate that the same is true here in the remainder of this work, which is extremely useful in our case because we aim to presume ignorance of what the best-fit parameters are when generating the analytical covariance matrix and the compression matrix.

### 05 Apr 2024 v3

@@ 4.1.1 Application to data
We test the procedure’s suitability described above by constructing individual analytical covariance matrices for each of the nine clustering samples we consider. We then treat the full analytical covariance matrix as a block diagonal matrix with the individual covariance matrix sitting in its corresponding block. We assume there is no correlation between different surveys mainly because of the large cosmological distances and weak overlap between the different samples. However within the BOSS survey, there is overlap between the $z1$ and $z2$ bins, which cannot be modelled without extending the analytical covariance matrix work of Wadekar et al. (2020). For simplicity, we first test the effect of neglecting these cross-correlation terms entirely instead.
We fit our sample with the full simulated covariance matrix and a block-diagonal simulated covariance matrix, where we remove the cross-correlation between the BOSS samples by replacing all non-block-diagonal terms with zero. The priors for our fit are summarized in Table 1, and match those in Glanville et al. (2022) where we fit the same data set with PyBird. The results are summarized in Fig 2 and Table 2. We found negligible shifts in the mean of the posteriors before and after removing the cross-correlation. There was a small change of <10% in the parameter uncertainties for $A_{s}$ and $h$, but this difference is tiny. It is also not clear if this is due to sampling noise, because of the absence of physical correlation between these two bins, or because the simulated cross-correlation itself is noisy. As such, we deem it reasonable to simply ignore this cross-correlation component in our fits such that we compare the analytical and simulated covariance matrices on an equal footing.
@@
Table 1: The priors on the cosmological parameters and the nuisance parameters for this work. We follow D'Amico et al. (2020) to set $c_{2}=\frac{b_{2}+b_{4}}{\sqrt{2}}$ and $c_{4}=\frac{b_{2}-b_{4}}{\sqrt{2}}=0$ because they found $b_{2}$ and $b_{4}$ are generally anti-correlated in simulations and observational data. Here, $\mathcal{U}$($\mathcal{N}$) denotes the uniform (Gaussian) prior where the first number is the lower bound (mean) and the second number is the upper bound (standard deviation). These priors are used and validated by Glanville et al. (2022) to fit the same data set. $A_{s}^{\mathrm{fid}},h^{\mathrm{fid}},$ and $(\Omega_{\mathrm{cdm}}h^{2})^{\mathrm{fid}}$ denotes the fiducial cosmological parameters used to calculate the center of the pre-computed grid of power spectrum. The widths of the priors (0.8, 0.08, and 0.04 for $A_{s}^{\mathrm{fid}},h^{\mathrm{fid}},$ and $(\Omega_{\mathrm{cdm}}h^{2})^{\mathrm{fid}}$ respectively) are to ensure cosmological parameters outside the grid won’t be evaluated. The fiducial parameter values are in Table 3, but they are changed depending on the exact tests we run.
Figure 2: This plot compares the constraints from the Full simulated covariance matrix with the ones from the block-diagonal simulated covariance matrix. Table 2 summarizes the constraints of the cosmological parameters. In general, we see no significant changes in the cosmological parameters, which indicates ignoring the cross-correlation does not affect the mean of the posterior for our data set.
@@
Figure 3: This plot compares the analytical covariance matrix with the simulated one for the eBOSS LRG NGC sample. The analytical covariance matrix is generated using the best-fit cosmological and bias parameters from the first iteration, which itself used the “Template" parameters in Table 3 and no MOPED compression (A1 in Table 4). The blue band corresponds to the sampling noise of the simulated covariance matrix. From left to right, we show the auto-covariance matrix for the monopole, the cross-covariance matrix for the monopole and quadrupole, and the auto-covariance matrix for the quadrupole. From top to bottom, we fix the first wavevector of the covariance matrix to $0.045h\rm{Mpc}^{-1}$, $0.095h\rm{Mpc}^{-1}$, $0.145h\rm{Mpc}^{-1}$, and $0.195h\rm{Mpc}^{-1}$ and vary the second wavevector from $0.0h\rm{Mpc}^{-1}$ to $0.30h\rm{Mpc}^{-1}$. The plots show that the analytical auto-covariance matrix is generally consistent with the simulated covariance matrix within the sampling noise limit, especially along the diagonal. The small disagreement along the off-diagonal does not affect the cosmological constraints as shown in section 5.
Fig. 3 compares the analytical covariance matrix with the simulated covariance matrix for the eBOSS LRG NGC subsample. For the analytical covariance matrix, we use the best-fit cosmological parameters obtained from our first fitting iteration which itself used an analytical covariance matrix computed with the “Template" parameters in Table 3, and without the compression (A1 in Table 4). [^4] You can find the comparison of the covariance matrix for the other eight samples here https://github.com/YanxiangL/Analytical-compressed-cov/tree/main/analytic_vs_sim. The blue band is the sampling noise of the simulated covariance matrix computed with equation (2). Fig 3 illustrates that the analytical covariance matrix is generally consistent with the simulated covariance matrix within the sampling noise, especially along the diagonal. The small disagreement could be because the local Lagrangian relation is not perfect, so some part of the analytical covariance matrix deviates slightly from the simulated one. Nonetheless, we find that the disagreement in the off-diagonal elements of the covariance matrix does not significantly change the constraints on the cosmological parameters, as the amplitude of these components is generally much smaller than either the diagonals of their respective blocks. Wadekar et al. (2020) similarly demonstrates that the off-diagonal components of the covariance between the monopole and quadrupole have a marginal impact on the constraint of the cosmological parameters.
@@ 4.2 The MOPED algorithm
The Massively Optimized Parameter Estimation and Data compression (MOPED) algorithm has been demonstrated to be able to reduce the number of data points without losing information (Heavens et al., 2000; Heavens et al., 2017). The algorithm achieves this by keeping the Fisher Matrix of the statistic in question with respect to the parameters of interest the same before and after the compression. Section 2 demonstrates the compressed precision matrix gives a smaller fractional error and bias than the uncompressed precision matrix. Therefore, the compression could also provide more reliable constraints for the cosmological parameters. We demonstrate its application to the EFTofLSS model and our large-scale structure dataset here.
@@
where the model power spectrum is given by $\boldsymbol{P}$ and $\boldsymbol{P}_{,m}$ is the partial derivative of the model power spectrum with respect to the $m^{\mathrm{th}}$ free parameter. The transpose is denoted by $T$ in equation (18).
To compute the compression matrix $\operatorname{\boldsymbol{\mathsf{D}}}$, we need to know the derivative of the model power spectrum with respect to the cosmological and nuisance parameters. For the derivatives with respect to the cosmological parameters, this is done when calculating the Taylor expansion of the power spectrum. A detailed description is given in section 4.3. In this section, we mainly focus on the derivatives of the power spectrum with respect to the nuisance parameters. In PyBird, the non-linear power spectrum calculation is separated into 24 different terms (3 linear terms, 12 loop terms, 6 counter terms, and 3 stochastic terms). Each of these terms is independent of the nuisance parameters. More importantly, the final non-linear power spectrum is the linear combination of these 24 different terms multiplied by the appropriate bias parameters
@@
The MOPED compression changes the covariance matrix and the power spectrum, so we must change the likelihood function accordingly. Before applying the compression, the power spectrum is expected to follow a Gaussian distribution, so the likelihood function $\mathcal{L}$ is given by
@@
where $\boldsymbol{P}_{d}$ is the data power spectrum and $\boldsymbol{P}^{W}=\operatorname{\boldsymbol{\mathsf{W}}}\boldsymbol{P_{l}}$ is the model power spectrum after convolving with the window function $\operatorname{\boldsymbol{\mathsf{W}}}$. In this work, both the analytical and simulated covariance matrix take the survey window function into account. Therefore, to calculate the compression matrix $\operatorname{\boldsymbol{\mathsf{D}}}$, we need to input the derivative of the non-linear power spectrum convolved with the window function ($\boldsymbol{P}^{W}$) with respect to the cosmological and nuisance parameters. Furthermore, this means the compressed windowed non-linear power spectrum is given by the compression matrix $\operatorname{\boldsymbol{\mathsf{D}}}$ multiplied by the windowed non-linear power spectrum ($\boldsymbol{P}^{W}$). From Heavens et al. (2000), the covariance matrix after compression is the identity matrix, so the likelihood after applying MOPED is given by
@@
To speed up the MCMC, PyBird analytically marginalizes over $\vec{b}_{\mathrm{NG}}$ (see equation (9)). The likelihood function after the analytical marginalization is given by (D'Amico et al., 2020)
@@
$F_{2,ij}=(P^{W}_{l,\mathrm{lin},i})^{T}\operatorname{\boldsymbol{\Psi}}P^{W}_{l,\mathrm{lin},j}$
$F_{1,i}=(P^{W}_{l,\mathrm{const}})^{T}\operatorname{\boldsymbol{\Psi}}P^{W}_{l,\mathrm{lin},i}+P_{d}^{T}\operatorname{\boldsymbol{\Psi}}P^{W}_{l,\mathrm{lin},i}$
$F_{0}=-\frac{1}{2}(P^{W}_{l,\mathrm{const}})^{T}\operatorname{\boldsymbol{\Psi}}P^{W}_{l,\mathrm{const}}+(P^{W}_{l,\mathrm{const}})^{T}\operatorname{\boldsymbol{\Psi}}P_{d}-\frac{1}{2}P_{d}^{T}\operatorname{\boldsymbol{\Psi}}P_{d},$ (23)
where $P^{W}_{l,\mathrm{lin},i}$ and $P^{W}_{l,\mathrm{const}}$ are $P_{l,\mathrm{lin},i}$ and $P_{l,\mathrm{const}}$ from equation (11) and (10) respectively after convolving with the window function.
@@
Now, we have the marginalized likelihood function. The next obstacle is applying the MOPED compression to $P^{W}_{l,\mathrm{lin},i}$ and $P^{W}_{l,\mathrm{const}}$. Equation (8) shows the windowed nonlinear power spectrum $\boldsymbol{P}^{W}$ is just a linear combination of $P^{W}_{l,\mathrm{lin},i}$ and $P^{W}_{l,\mathrm{const}}$. The bias parameters are just constants and would not affect the matrix multiplication, so we have
$DP^{W}=\sum_{i}P^{W}_{l,\mathrm{lin},i}+DP^{W}_{l,\mathrm{const}}.$ (24)
@@
This means we can calculate the compression matrix for $P^{W}_{l,\mathrm{lin},i}$ and $P^{W}_{l,\mathrm{const}}$ the same way as $P^{W}$, by multiplying the compression matrix to it. After some mathematical manipulation, the expression for the marginalized likelihood stays the same, but one needs to change equation (23) to
$F_{2,ij}^{\mathrm{comp}}=(DP^{W}_{l,\mathrm{lin},i})^{T}DP^{W}_{l,\mathrm{lin},j}$
$F_{1,i}^{\mathrm{comp}}=(DP^{W}_{l,\mathrm{lin},i})^{T}DP^{W}_{l,\mathrm{const}}+(DP^{W}_{l,\mathrm{lin},i})^{T}DP_{d}$
$F_{0}^{\mathrm{comp}}=-\frac{1}{2}(DP^{W}_{l,\mathrm{const}})^{T}DP^{W}_{l,\mathrm{const}}+(DP_{d})^{T}DP^{W}_{l,\mathrm{const}}-$
$\frac{1}{2}(DP_{d})^{T}DP_{d}.$ (25)
@@
There are many ways to calculate the equation (25) by rearranging the order of the matrix/vector multiplications. For example, for $F_{2,ij}^{\mathrm{comp}}$ we can write it as $(P^{W}_{l,\mathrm{lin},i})^{T}D^{T}DP^{W}_{l,\mathrm{lin},j}$. However, it is not computationally efficient to calculate $F_{2,ij}^{\mathrm{comp}}$ this way because $D^{T}D$ has the same dimension as the uncompressed precision matrix, which will not speed up the code. Having investigated different approaches, we found the most computationally efficient way was to pre-compute $DP_{d}$ and $(DP_{d})^{T}DP_{d}$ because the data vector does not change during MCMC. Then during each step in MCMC, we calculate $DP^{W}_{l,\mathrm{lin},i}$ and $DP^{W}_{l,\mathrm{const}}$ first, then use the transpose operation to find $(DP^{W}_{l,\mathrm{lin},i})^{T}$ and $(DP^{W}_{l,\mathrm{const}})^{T}$ because the transpose operation is much faster than matrix multiplication. After transposing, both the model and data power spectrum are compressed, so the matrix multiplication is faster. We then substitute the compressed power spectra into equation (25) and equation (22) to find the marginalized likelihood.
@@
Compression has several benefits. The MOPED algorithm significantly reduces the dimension of the data vector. Therefore, we will need fewer simulations to estimate the covariance matrix to reach the same precision as shown in Fig 1. Furthermore, the chi-squared calculation is one of the slowest parts when sampling the posterior with MCMC. Reducing the size of the data vector speeds up the MCMC algorithm. Lastly, we don’t have to multiply by the covariance matrix after the compression, which also speeds up the MCMC.
@@
However, there are also caveats — we need to multiply the compression matrix with the model power spectrum at every step. This could cancel out the potential speed-up. The second caveat is that theoretically, the compression matrix needs to be generated with the best-fit parameters for the compression to be lossless. Heavens et al. (2000) found that using parameters that are different from the best-fit only marginally increases the credible regions. We demonstrate that the same is true here in the remainder of this work. This is extremely useful in our case because we aim to presume ignorance of the best-fit parameters when generating the analytical covariance matrix and the compression matrix.